# Informativity Conditions for Multiple Signals: Properties, Experimental Design, and Applications (extended version)

### v3

@@ Abstract
Recent studies highlight the importance of persistently exciting condition in single signal sequence for model identification and data-driven control techniques. However, maintaining prolonged excitation in control signals introduces significant challenges, as continuous excitation can reduce the lifetime of mechanical devices. In this paper, we introduce three informativity conditions for various types of multi-signal data, each augmented by weight factors. We explore the interrelations between these conditions and their rank properties in linear time-invariant systems. All three conditions can extend Willems’ fundamental lemma and are utilized to assess the properties of the system. Furthermore, we introduce open-loop experimental design methods tailored to each of the three conditions, which can synthesize the required excitation conditions either offline or online, even in the presence of limited information within each signal segment. Illustrative examples confirm that these conditions yield satisfactory outcomes in both least-squares identification and the construction of data-driven controllers.
@@ 1 Introduction
With the continuous development and application of data-driven technologies, the academic interest in data informativity has surged notably. In this context, the concept of persistently exciting (PE) condition (Willems et al., 2005) plays a crucial role. The PE condition requires that a signal be sufficiently rich over time to capture all dynamic characteristics of the system. This requirement is fundamental to various system identification methods (Ioannou and Fidan, 2006; Katayama et al., 2005; Markovsky et al., 2005; Markovsky, 2012; Lovera et al., 2000) and data-driven control methods (Berberich et al., 2021; Bongard et al., 2023; Liu et al., 2023; De Persis and Tesi, 2020; Park and Ikeda, 2009; Nortmann and Mylvaganam, 2023).
@@
While persistency of excitation can indeed be enforced in theory by well-known inputs (Gevers et al., 2009) such as pseudo-random binary sequences (PRBS), broadband white noise, or periodic signals with many harmonics, their practical deployment is often constrained. First, guaranteeing PE with these signals typically requires a long single trajectory, which may not be available in experimental environments where data collection must be interrupted or restarted. Second, long-duration stochastic or high-frequency excitations can violate safety or performance constraints in many systems, e.g., process reactors or power electronic converters where sustained probing may be unsafe (Yang et al., 2024), or robotic platforms where prolonged excitation induces mechanical wear (Hamaya et al., 2021). Third, when multiple shorter trajectories are collected, naive juxtaposition of PRBS-type signals does not necessarily yield a full-rank Hankel matrix due to phase misalignment or heterogeneous signal magnitudes. These limitations highlight the need for conditions that certify data informativity beyond the classical single-trajectory PE framework.
@@
Existing research has shown that PE condition is not a necessary prerequisite for system identification and data-driven control design (van Waarde et al., 2020b; Kang and You, 2023; Van Waarde et al., 2023). In the domain of system identification, methods such as dynamic regressor extension and mixing (Aranovskiy et al., 2016; Ortega et al., 2021) have relaxed the PE condition by extending the regression model in the time domain. Certain adaptive methods (Dhar et al., 2022) only require that excitation conditions be met during the initial phase. Furthermore, the challenge of insufficient information from a single sensor has been addressed through distributed frameworks (Chen et al., 2014; Xie and Guo, 2018) and centralized frameworks (Markovsky and Pintelon, 2015). Despite these advancements, these methods often relax the information content of system regression vectors or system trajectories, yet rarely focus on designing from the perspective of control signals. Therefore, a clear experimental design strategy for system input signals remains essential. While current experimental design theories are primarily based on the premise of satisfying identifiability (De Persis and Tesi, 2021; van Waarde, 2022; Bombois et al., 2006; Hjalmarsson, 2005), there remains a significant gap in research regarding how to design experiments that guarantee the effectiveness of identification methods when the informativeness of input signals is lacking. In such cases, the use of multiple trajectories may be necessary.
@@
In the design of data-driven control methods, A new control framework is developed in (van Waarde et al., 2020b) that eliminates the stringent requirement of PE, but the trajectory data used still needs to guarantee the identifiability condition. The collectively persistently exciting (CPE) condition proposed in (van Waarde et al., 2020a) and (Yu et al., 2021) offers potential insights for the design of data-driven control methods using trajectories generated by insufficiently informative controllers. By organizing the Hankel matrices that evaluate the excitability of each signal sequence side-by-side into a larger rank judgment matrix, this condition extends the fundamental lemma from a single signal sequence to multiple signal sequences. However, this large matrix structure can lead to a dramatic increase in computational burden as the number of columns increases, and excessive differences in the magnitude of the values between each Hankel matrix can also cause deviations from the expected results. Additionally, it is currently an open question how to provide more choices of matrix structures for different data-driven problems (Markovsky and Dörfler, 2021).
@@
The remainder of this paper is organized as follows. Section 2 introduces the preliminaries, covering key definitions and a review of prior work on the PE condition. In Section 3, we define three CPE conditions tied to data dimensions, analyze their transformational relationships and rank properties, and extend Willems’ fundamental lemma. Section 4 introduces experimental methods tailored to the three CPE conditions and demonstrates their validity. Section 5 introduces an identifier for single system, extending the results to the design of distributed adaptive identifiers with exponential convergence rates under CPE conditions. Section 6 provides illustrative examples of applications to validate the methodologies and theories discussed, while Section 7 concludes the paper.
@@ 2.2 Persistency of Excitation for a Single Sequence
The concept of persistency of excitation for a single signal sequence, as defined in Willems et al. (2005), is elucidated below.
@@
(Persistently Exciting Condition (Willems et al., 2005)). A signal sequence $z^{[0,T_{0}-1]}$ with $z(k)\in\mathbb{R}^{m}$, $k=0,...,T_{0}-1$ is PE of order $L$ if the block Hankel matrix $H_{L}(z^{[0,T_{0}-1]})$ has full row rank $mL$.
@@
where $A\in{\mathbb{R}^{n\times n}}$, ${B}\in{\mathbb{R}^{n\times{m}}}$, ${x(k)}\in{\mathbb{R}^{n}}$ and ${u(k)}\in{\mathbb{R}^{m}}$ represent the state vector and control input vector, respectively. The pair $(A,B)$ is controllable. A fundamental property in (Willems et al., 2005, Cor. 2) demonstrates that the LTI system (1) satisfies the rank condition
@@
This property indicates that the image of the matrix $\begin{bmatrix}{{H_{L}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{L}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}$ is the representation of the $L$-long state-input trajectory space of the LTI system (1). It has been rigorously proved in the behavioral system framework (Willems et al., 2005, Th. 1) and the state-space framework (De Persis and Tesi, 2020, Lem. 2), respectively.
@@
Indeed, if the input sequence $u^{[0,{T_{0}}-1]}$ of system (1) is PE of order $n+1$, the matrix $\begin{bmatrix}{{H_{1}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{1}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}$ has full row rank. This condition enables the recovery of the system matrices $A$ and $B$ from the data $x^{[0,{T_{0}}-1]}$ and $u^{[0,{T_{0}}-1]}$, i.e., identifiability condition (see (Markovsky and Dörfler, 2023; van Waarde et al., 2020b)). The system matrices can be uniquely determined by solving the following equation:
@@
On the other hand, under the influence of the prior control input $u^{[0,T_{0}-1]}$ that satisfies the PE condition, the system trajectories ${\bar{x}}^{[0,L-1]}$ and ${\bar{u}}^{[0,L-1]}$ can form the non-parametric representation in (3), which solely relies on the prior data. Consequently, in addition to its important role in system identification, PE condition has received widespread attention in the development of data-driven controllers such as parametric model reconstruction methods (De Persis and Tesi, 2020; Nortmann and Mylvaganam, 2023) and data-driven model predictive control methods (Berberich et al., 2021; Bongard et al., 2023; Liu et al., 2023).
Even though it has been shown how to ensure the informativeness of a single control sequence (Gevers et al., 2009), in some cases it may still not be possible to generate sufficiently informative trajectories due to the many limitations of the experiment. However, designing experiments with insufficiently informative controllers fail to ensure the identifiability condition and the feasibility of data-driven approaches.
@@ 3.1 Three Collectively Persistently Exciting Conditions
The MCPE condition does not specify the number of dimensions of each signal sequence, i.e., the $T_{i}$ of each signal can be different. Only the sum of all signal lengths is required to satisfy $\sum\limits_{i=1}^{p}{{T_{i}}}\geq mL+p(L-1)$. However, this flexibility results in an unrestricted number of columns in the rank judgment matrix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$}, potentially leading to an excessive number of redundant columns in the matrix. This abundance of columns can significantly increase computational complexity in practical applications such as model predictive control Berberich et al. (2021); Bongard et al. (2023); Liu et al. (2023).
@@
Next, we introduce a hybrid collectively persistently exciting (HCPE) condition for partially same-dimensional signal sequences. This condition employs a rank judgment matrix that combines the matrix structures from both the CCPE condition and the MCPE condition, leveraging the advantages and mitigating the disadvantages of each.
@@
The benefit of MCPE lies not in reducing the number of matrix columns, but in simplifying the overall computational procedure of informativity verification while enhancing flexibility in handling heterogeneous trajectories. Instead of repeatedly checking PE on each trajectory or artificially aligning all trajectories to the same horizon (as required by CCPE), MCPE organizes all heterogeneous trajectories into a single rank test. Thus, only one unified check is needed, while weighting factors mitigate numerical ill-conditioning caused by magnitude discrepancies across datasets. Although the resulting matrix may be wider, the computation pipeline is considerably simpler in certain scenarios, especially when handling heterogeneous-length data, which is shown with the illustrative example in the Section 6.5.
@@
The CPE condition proposed in van Waarde et al. (2020a) shares similarities with the first type introduced in this paper, the MCPE condition, as both combine Hankel matrices in a side-by-side fashion. However, the MCPE condition distinguishes itself by incorporating weight factors, $\alpha_{i}$, which allows it to perform better when handling data with significant order-of-magnitude differences. Additionally, a careful selection of the weighting factors can enhance both performance and computational efficiency when employing data-driven methods, as demonstrated in the examples in Section 6.2. Notably, despite the introduction of weight factors, the difficulty of synthesis remains unchanged. This is attributed to the fact that for any $\alpha_{i}\neq 0$, the rank of the matrix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$} remains the same as when $\alpha_{i}=1$.
@@ 3.2 Transformation Relations and rank properties of CPE Conditions
      $T_{0}=T_{1}=...=T_{p}$, $\text{im}\,{{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{T}}\cap\text{leftker}\,{\textbf{1}_{p}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE;
@@
      $T_{0}=T_{1}=...=T_{\bar{p}}$, $\text{im}\,{{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{\bar{p}})}^{T}}\cap\text{leftker}\,{\begin{bmatrix}{{\textbf{1}_{\bar{p}}}}&{{0_{\bar{p}\times(p-\bar{p})}}}\\ {{0_{p-\bar{p}}}}&{{I_{p-\bar{p}}}}\end{bmatrix}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also HCPE;
@@
      $T_{0}=T_{\bar{p}+1}=...=T_{p}$, $\text{im}\,{{H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{T}}\cap\text{leftker}\,{\textbf{1}_{\bar{p}+1}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE.
@@ 3.3 Extension of Fundamental Lemma
${{\mathcal{O}}_{L}}={\begin{bmatrix}I&{{A^{T}}}&{{{({A^{2}})}^{T}}}&\cdots&{{{({A^{L-1}})}^{T}}}\end{bmatrix}^{T}}.$
@@
By introducing CCPE, MCPE, and HCPE, the extended lemma formalizes the conditions under which segmented data collectively provide sufficient information for trajectory representation. This generalization broadens the lemma’s applicability to scenarios such as systems with distributed sensing, segmented experiments, or partial data. Notably, the work in van Waarde et al. (2020a) marks an initial step toward addressing multiple trajectories in Willems’ Lemma, while Lemma 2 strictly extends and incorporates the results from van Waarde et al. (2020a). In special or complex cases (e.g., significant numerical discrepancies between trajectories or equal trajectory lengths), the additional flexibility afforded by Lemma 2 enhances its applicability.
@@
Willems’ fundamental lemma has become a cornerstone in the development of data-driven control methodologies. Lemmas 1 and 2 share core properties with Williams’ lemma, positioning them as fundamental tools for similar applications. Specifically, Lemma 1 enables results analogous to (De Persis and Tesi, 2020, Th. 3), where the problem of stabilizing controller design is reformulated as a linear matrix inequality (LMI) optimization task. Additionally, Lemma 2 supports the construction of predictive models, providing a foundation for data-driven model predictive control (MPC) designs, as demonstrated in works such as Berberich et al. (2021); Bongard et al. (2023); Liu et al. (2023). Therefore, Lemmas 1 and 2 extend the applicability of data-driven techniques to scenarios involving multiple a priori trajectories, offering broader prospects for practical applications. Moreover, while the identification methods and extension lemma presented in this paper are based on fully observable LTI systems, they can be easily extended to input-output state-space systems, as the core application of these methods is rooted in Lemma 1 , which holds for input-output systems as well.
@@
This condition is analogous to the informativeness condition in adaptive methods. According to Theorem 1, the same holds for both the CCPE and HCPE conditions. Consequently, the informativeness condition defined in this paper has significant potential for application to adaptive data-driven methods as well. Detailed results regarding the three conditions in distributed adaptive system identification methods are elaborated in the Section 5.
4 Experimental Design for CPE Conditions

### v4

@@ Abstract
Recent studies highlight the importance of the persistently exciting condition in a single signal sequence for model identification and data-driven control techniques. However, maintaining prolonged excitation in control signals introduces significant challenges, as continuous excitation can reduce the lifetime of mechanical devices. In this paper, we introduce three informativity conditions for various types of multi-signal data, each augmented with weight factors. We explore the interrelations between these conditions and their rank properties in linear time-invariant systems. All three conditions can extend Willems’ fundamental lemma and are utilized to assess the properties of the system. Furthermore, we introduce open-loop experimental design methods tailored to each of the three conditions, which can synthesize the required excitation conditions either offline or online, even in the presence of limited information within each signal segment. Illustrative examples confirm that these conditions yield satisfactory outcomes in both least-squares identification and the construction of data-driven controllers.
@@ 1 Introduction
With the continuous development and application of data-driven technologies, the academic interest in data informativity has surged notably. In this context, the concept of persistently exciting (PE) condition (WILLEMS2005325) plays a crucial role. The PE condition requires that a signal be sufficiently rich over time to capture all dynamic characteristics of the system. This requirement is fundamental to various system identification methods (ioannou2006adaptive; katayama2005subspace; MARKOVSKY2005755; markovsky2012low; LOVERA20001639) and data-driven control methods (9109670; 9744574; 9903320; 8933093; PARK20091265; 10124991).
@@
While persistency of excitation can indeed be enforced in theory by well-known inputs (5325719) such as pseudo-random binary sequences (PRBS), broadband white noise, or periodic signals with many harmonics, their practical deployment is often constrained. First, guaranteeing PE with these signals typically requires a long single trajectory, which may not be available in experimental environments where data collection must be interrupted or restarted. Second, long-duration stochastic or high-frequency excitations can violate safety or performance constraints in many systems, e.g., process reactors or power electronic converters where sustained probing may be unsafe (YANG2024100986), or robotic platforms where prolonged excitation induces mechanical wear (doi:10.1177/0278364919853618). Third, when multiple shorter trajectories are collected, naive juxtaposition of PRBS-type signals does not necessarily yield a full-rank Hankel matrix due to phase misalignment or heterogeneous signal magnitudes. These limitations highlight the need for conditions that certify data informativity beyond the classical single-trajectory PE framework.
@@
Existing research has shown that the PE condition is not a necessary prerequisite for system identification and data-driven control design (8960476; KANG2023111130; 10317631). In the domain of system identification, methods such as dynamic regressor extension and mixing (7526771; 9121700) have relaxed the PE condition by extending the regression model in the time domain. Certain adaptive methods (DHAR2022100672) only require that excitation conditions be met during the initial phase. Furthermore, the challenge of insufficient information from a single sensor has been addressed through distributed frameworks (W6578135; doi:10.1137/16M1106791) and centralized frameworks (7098430). Despite these advancements, these methods often relax the information content of system regression vectors or system trajectories, yet rarely focus on designing from the perspective of control signals. Therefore, a clear experimental design strategy for system input signals remains essential. While current experimental design theories are primarily based on the premise of satisfying identifiability (DEPERSIS2021285; 9406124; BOMBOIS20061651; HJALMARSSON2005393), there remains a significant gap in research regarding how to design experiments that guarantee the effectiveness of identification methods when the informativeness of input signals is lacking. In such cases, the use of multiple trajectories may be necessary.
@@
In the design of data-driven control methods, a new control framework is developed in (8960476) that eliminates the stringent requirement of PE, but the trajectory data used still needs to guarantee the identifiability condition. The collectively persistently exciting (CPE) condition proposed in (9062331) and (9682952) offers potential insights for the design of data-driven control methods using trajectories generated by insufficiently informative controllers. By organizing the Hankel matrices that evaluate the excitability of each signal sequence side-by-side into a larger rank judgment matrix, this condition extends the fundamental lemma from a single signal sequence to multiple signal sequences. However, this large matrix structure can lead to a dramatic increase in computational burden as the number of columns increases, and excessive differences in the magnitude of the values between each Hankel matrix can also cause deviations from the expected results. Additionally, it is currently an open question how to provide more choices of matrix structures for different data-driven problems (MARKOVSKY202142).
@@
The remainder of this paper is organized as follows. Section 2 introduces the preliminaries, covering key definitions and a review of prior work on the PE condition. In Section 3, we define three CPE conditions tied to data dimensions, analyze their transformational relationships and rank properties, and extend Willems’ fundamental lemma. Section 4 introduces experimental methods tailored to the three CPE conditions and demonstrates their validity. Section 6 provides illustrative examples of applications to validate the methodologies and theories discussed, while Section 7 concludes the paper.
@@ 2.2 Persistency of Excitation for a Single Sequence
The concept of persistency of excitation for a single signal sequence, as defined in WILLEMS2005325, is elucidated below.
@@
( (WILLEMS2005325)). A signal sequence $z^{[0,T_{0}-1]}$ with $z(k)\in\mathbb{R}^{m}$, $k=0,...,T_{0}-1$ is PE of order $L$ if the block Hankel matrix $H_{L}(z^{[0,T_{0}-1]})$ has full row rank $mL$.
@@
where $A\in{\mathbb{R}^{n\times n}}$, ${B}\in{\mathbb{R}^{n\times{m}}}$, ${x(k)}\in{\mathbb{R}^{n}}$ and ${u(k)}\in{\mathbb{R}^{m}}$ represent the state vector and control input vector, respectively. The pair $(A,B)$ is controllable. A fundamental property in (WILLEMS2005325, Cor. 2) demonstrates that the LTI system (1) satisfies the rank condition
@@
This property indicates that the image of the matrix $\begin{bmatrix}{{H_{L}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{L}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}$ is the representation of the $L$-long state-input trajectory space of the LTI system (1). It has been rigorously proved in the behavioral system framework (WILLEMS2005325, Th. 1) and the state-space framework (8933093, Lem. 2), respectively.
@@
Indeed, if the input sequence $u^{[0,{T_{0}}-1]}$ of system (1) is PE of order $n+1$, the matrix $\begin{bmatrix}{{H_{1}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{1}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}$ has full row rank. This condition enables the recovery of the system matrices $A$ and $B$ from the data $x^{[0,{T_{0}}-1]}$ and $u^{[0,{T_{0}}-1]}$, i.e., identifiability condition (see (9904308; 8960476)). The system matrices can be uniquely determined by solving the following equation:
@@
On the other hand, under the influence of the prior control input $u^{[0,T_{0}-1]}$ that satisfies the PE condition, the system trajectories ${\bar{x}}^{[0,L-1]}$ and ${\bar{u}}^{[0,L-1]}$ can form the non-parametric representation in (3), which solely relies on the prior data. Consequently, in addition to its important role in system identification, the PE condition has received widespread attention in the development of data-driven controllers such as parametric model reconstruction methods (8933093; 10124991) and data-driven model predictive control methods (9109670; 9744574; 9903320).
Even though it has been shown how to ensure the informativeness of a single control sequence (5325719), in some cases it may still not be possible to generate sufficiently informative trajectories due to the many limitations of the experiment. However, designing experiments with insufficiently informative controllers fails to ensure the identifiability condition and the feasibility of data-driven approaches.
@@ 3.1 Three Collectively Persistently Exciting Conditions
The MCPE condition does not specify the number of dimensions of each signal sequence, i.e., the $T_{i}$ of each signal can be different. Only the sum of all signal lengths is required to satisfy $\sum\limits_{i=1}^{p}{{T_{i}}}\geq mL+p(L-1)$. However, this flexibility results in an unrestricted number of columns in the rank test matrix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$}, potentially leading to an excessive number of redundant columns in the matrix. This abundance of columns can significantly increase computational complexity in practical applications such as model predictive control 9109670; 9744574; 9903320.
@@
Next, we introduce a hybrid collectively persistently exciting (HCPE) condition for partially same-dimensional signal sequences. This condition employs a rank test matrix that combines the matrix structures from both the CCPE condition and the MCPE condition, leveraging the advantages and mitigating the disadvantages of each.
@@
The benefit of MCPE lies not in reducing the number of matrix columns, but in simplifying the overall computational procedure of informativity verification while enhancing flexibility in handling heterogeneous trajectories. Instead of repeatedly checking PE on each trajectory or artificially aligning all trajectories to the same horizon (as required by CCPE), MCPE organizes all heterogeneous trajectories into a single rank test. Thus, only one unified check is needed, while weighting factors mitigate numerical ill-conditioning caused by magnitude discrepancies across datasets. Although the resulting matrix may be wider, the computation pipeline is considerably simpler in certain scenarios, especially when handling heterogeneous-length data, which is shown with the illustrative example in the Section 6.5. The CPE condition proposed in 9062331 shares similarities with the first type introduced in this paper, the MCPE condition, as both combine Hankel matrices in a side-by-side fashion. However, the MCPE condition distinguishes itself by incorporating weight factors, $\alpha_{i}$, which allows it to perform better when handling data with significant order-of-magnitude differences. Additionally, a careful selection of the weighting factors can enhance both performance and computational efficiency when employing data-driven methods, as demonstrated in the examples in Section 6.2. Notably, despite the introduction of weight factors, the difficulty of synthesis remains unchanged. This is attributed to the fact that for any $\alpha_{i}\neq 0$, the rank of the matrix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$} remains the same as when $\alpha_{i}=1$.
@@ 3.2 Transformation Relations and rank properties of CPE Conditions
      $T_{0}=T_{1}=...=T_{p}$, $\text{im}\,{{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{\top}}\cap\text{leftker}\,{\textbf{1}_{p}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE;
@@
      $T_{0}=T_{1}=...=T_{\bar{p}}$, $\text{im}\,{{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{\bar{p}})}^{\top}}\cap\text{leftker}\,{\begin{bmatrix}{{\textbf{1}_{\bar{p}}}}&{{0_{\bar{p}\times(p-\bar{p})}}}\\ {{0_{p-\bar{p}}}}&{{I_{p-\bar{p}}}}\end{bmatrix}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also HCPE;
@@
      $T_{0}=T_{\bar{p}+1}=...=T_{p}$, $\text{im}\,{{H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{\top}}\cap\text{leftker}\,{\textbf{1}_{\bar{p}+1}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE.
@@ 3.3 Extension of Fundamental Lemma
${{\mathcal{O}}_{L}}={\begin{bmatrix}I&{{A^{\top}}}&{{{({A^{2}})}^{\top}}}&\cdots&{{{({A^{L-1}})}^{\top}}}\end{bmatrix}^{\top}}$.
@@
By introducing CCPE, MCPE, and HCPE, the extended lemma formalizes the conditions under which segmented data collectively provide sufficient information for trajectory representation. This generalization broadens the lemma’s applicability to scenarios such as systems with distributed sensing, segmented experiments, or partial data. Notably, the work in 9062331 marks an initial step toward addressing multiple trajectories in Willems’ Lemma, while Lemma 2 strictly extends and incorporates the results from 9062331. In special or complex cases (e.g., significant numerical discrepancies between trajectories or equal trajectory lengths), the additional flexibility afforded by Lemma 2 enhances its applicability.
@@
Willems’ fundamental lemma has become a cornerstone in the development of data-driven control methodologies. Lemmas 1 and 2 share core properties with Willems’ lemma, positioning them as fundamental tools for similar applications. Specifically, Lemma 1 enables results analogous to (8933093, Th. 3), where the problem of stabilizing controller design is reformulated as a linear matrix inequality (LMI) optimization task. Additionally, Lemma 2 supports the construction of predictive models, providing a foundation for data-driven model predictive control (MPC) designs, as demonstrated in works such as 9109670; 9744574; 9903320. Therefore, Lemmas 1 and 2 extend the applicability of data-driven techniques to scenarios involving multiple a priori trajectories, offering broader prospects for practical applications. Moreover, while the identification methods and extension lemma presented in this paper are based on fully observable LTI systems, they can be easily extended to input-output state-space systems, as the core application of these methods is rooted in Lemma 1 , which holds for input-output systems as well.
@@
This condition is analogous to the informativeness condition in adaptive methods. According to Theorem 1, the same holds for both the CCPE and HCPE conditions. Consequently, the informativeness condition defined in this paper has significant potential for application to adaptive data-driven methods as well. Detailed results regarding the three conditions in distributed adaptive system identification methods are elaborated in the extended version (cao2025informativityconditionsmultiplesignals).
4 Experimental Design for the CPE Conditions